OpenStudy (anonymous):

√(2x+4)-√(x)=2 I know x=16, but how does one find this solution?

OpenStudy (mathstudent55):

Isolate one root, preferably sqrt(2x + 4), and square both sides. Simplify everything, then isolate the root again and square both sides again. Then solve the equation. At the end, check every solution in the original equation because squaring both sides can introduce extraneous solutions.

OpenStudy (anonymous):

I am confused. Isolate one root: root(3x+4)=root(x)+2 square bother sides: 3x+4=x+4 Isolate what root? Both roots have been squared. If I simplify it from here all I get is 0.

OpenStudy (mathstudent55):

You're on the right track, but you need to be careful when you square a sum. (x + y)^2 =/= x^2 + y^2

OpenStudy (mathstudent55):